OpenStudy (anonymous):

Part 1: Using complete sentences, explain how to find the equation of the line, in standard form and slope–intercept form, passing through (3, 6) and (–2, –4). (4 points) Part 2: Compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form. (3 points)

OpenStudy (anonymous):

Step 1 : Find the slope of the given points using m = y2 - y1/x2 - x1 Step 2: Plugin the values of x1, y1 and m in y-y1 = m (x - x1)